Plainwell Polymers is going to producing the main dashboard structure for a new car design. The dashboard will be produced in three colors with expected daily demand as follows:

Black: 530 units/day

Grey: 135 units/day

Tan: 185 units/day

The injection-molding machine can produce a total of 1,010 parts per day running 24 hours per day.

It costs Plainwell $1,110 to set up each color for producition and the daily holding cost for each part is $0.03. What is the length of the lowest cost cyclic schedule (ignoring setup time constraints) in days?
Answer #1

We have given following data Daily demand = d = (530+ 135+185) cmits/day d = 850 units/day so that Annual demand = = 850*365